OpenStudy (anonymous):

If 3^17+3^17+3^17+3^17+3^17+3^17+3^17+3^17+3^17+3^17=81^N, then th value of N is: a)6 b)17/4 c)17/2 d)19/4 e)16/9

OpenStudy (anonymous):

i think it is b

OpenStudy (anonymous):

can you please show some steps?

jimthompson5910 (jim_thompson5910):

If there are 9 copies of 3^17 being added then... 3^17+3^17+3^17+3^17+3^17+3^17+3^17+3^17+3^17=81^N 9*3^17=81^N (3^2)*(3^17)=(3^4)^N 3^(2 + 17) = 3^(4N) I'll let you finish

OpenStudy (anonymous):

There are 10 of 3^17

jimthompson5910 (jim_thompson5910):

Then none of the answer choices match

jimthompson5910 (jim_thompson5910):

If there are 9 of them added, then one of the answer choices matches.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if there are 9 of them the answer will be d since 4N=19, N=19/4

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

yeah I got it then, thank you so much for explaining this to me!! :)

jimthompson5910 (jim_thompson5910):

np